Consider the set of even single-digit number {0, 2, 4, 6, 8}.

·       Make a list of possible sample size of 2 that can be taken from this sets of numbers

·       Construct the sampling distribution of the sample means for the size of 3 and the standard error.



1
Expert's answer
2022-03-24T03:03:16-0400

Possible samples of size 2.

(0,2), (0,4), (0,6), (0,8), (2,4), (2,6), (2,8), (4,6), (4,8), (6,8)


Sampling distribution

Possible samples of size 3,

(0,2,4), (0,2,6), (0,2,8), (0,4,6), (0,4,8), (0,6,8), (2,4,6), (2,4,8), (2,6,8), (4,6,8)


"P(X=2)=P(X=2.67)=P(X=5.33)=P(X=6)=0.1"

"P(X=3.33)=P(X=4)=P(X=4.67)=0.2"


Standard Error




"\mu=∑xf(x)=4"

"\sigma=\sqrt{∑x^2f(x)-4^2}=\sqrt{17.33-4^2}"

"=\sigma^2=1.33"

"S.E=\frac{\sigma}{\sqrt{n}}=\frac{\sqrt{1.33}}{\sqrt{3}}=0.665"